OpenStudy (anonymous):

Monkey - Donkey=1 Elephant-Horse= ?

OpenStudy (anonymous):

3:P

OpenStudy (amistre64):

staten island zoo?

OpenStudy (anonymous):

lol how did 3 came can u plz explain?

OpenStudy (anonymous):

no its an iq question amistere

OpenStudy (anonymous):

not sure, i just guessed it

OpenStudy (amistre64):

i fail those rochester ink blots too ;)

OpenStudy (anonymous):

monkey n donkey have got one different alphabet

OpenStudy (anonymous):

amistre ........ur issues resolved with the admin;)

OpenStudy (anonymous):

wtf....

OpenStudy (dumbcow):

where do you get these Q's??

OpenStudy (anonymous):

monkey? donkey? ._.

OpenStudy (anonymous):

monkey - donkey = m -d .-.?

OpenStudy (anonymous):

what's the question lol >_<

OpenStudy (anonymous):

this is an iq question i got this in my medical entrance tezst.......

OpenStudy (anonymous):

people are posting funny questions lol, I am ENJOYING THIS! =D

OpenStudy (anonymous):

lmafo this is not funny....lol....its a serious question..lol

OpenStudy (anonymous):

donkey - monkey = 1? now tell me, how on Earth is that true LOL?

OpenStudy (anonymous):

i dont know how it became so am here to get help out of u..

OpenStudy (anonymous):

my final test is coming soon so i must practise asmuch as i have found the questions

OpenStudy (dumbcow):

oh ok that makes sense, its testing your ability to reason, recognize patterns, solve riddles...not really a math question

OpenStudy (anonymous):

hmmmm, elephant - horse = nothing since they have no letters in common ^_^

OpenStudy (dumbcow):

my guess is 3 , 5 or 8 because they only have the e and h in common, take those letters away and that leaves 5 distinct letters for elephant and 3 for horse

OpenStudy (anonymous):

I was half right~ lol

OpenStudy (anonymous):

it can't be 8 though

OpenStudy (anonymous):

hey guys why dont u try next question that have posted above..?

OpenStudy (anonymous):

the answer is 3 ^_^, I guess since there are 8 letters in elephant so (8) and 5 letters in horse - (5) = 3

OpenStudy (anonymous):

maybe .-.

OpenStudy (anonymous):

the way i see it is monkey / donkey is 6/6 = 1 so elephant/horse is 8/5=1 3/5